Suppose, the ADC in Atmega32 is configured to use 4 Volt as its voltage reference. Then Find: 

i. Step size, and

ii. Digital output in decimal, if Vin= 3 V.


1
Expert's answer
2021-01-31T09:03:10-0500

Controller ADC has 1023 steps and

i)   Step size with Vref=4V :         Step_size = 4/1023 = 3.91 mV. 


Digital data output will be Dout = Vin/step size.

ii)     Dout = 3V/(3.91mV)*1000 = 767